2. Suppose that the width of a rectangle is three feet shorter than length and that the perimeter of the rectangle is 84 feet.
a) Set up an equation for the perimeter involving only L, the length of the rectangle.
Answer:
b) Solve this equation algebraically to find the length of the rectangle. Find the width as well.
Answer: Length 22.5, Width 19.5
Show your work here:W=L-3
2W+2L=84
2(L-3)+2L=84
2L-6+2L=84
4L=84+6
4L=90
L=90/4
L=22.5 ANS. FOR THE LENGTH.
W=22.5-3=19.5 ANS. FOR THE WIDTH.
PROOF:
2*19.5+2*22.5=84
39+45=84
84=84

Answer by checkley77(12844)

W=L-3
2W+2L=84
2(L-3)+2L=84
2L-6+2L=84
4L=84+6
4L=90
L=90/4
L=22.5 ANS. FOR THE LENGTH.
W=22.5-3=19.5 ANS. FOR THE WIDTH.
PROOF:
2*19.5+2*22.5=84
39+45=84
84=84